### v2

@@ 4.5 Verification of Seamless Plugin Integration
DSCnormal-$\uparrow$ HD95normal-$\downarrow$ DSCnormal-$\uparrow$ HD95normal-$\downarrow$ DSCnormal-$\uparrow$ HD95normal-$\downarrow$
@@
4.6 Analysis on the Impact of Annotation Sparsity
To explore the impact of the annotation sparsity level over ProCNS and inspired by [7], we report the results of ProCNS obtained from training samples with five different sparse-versus-full annotated proportions. As shown in Table 6, ProCNS demonstrates superior performance across all five proportions, all of which surpass the performance of $\text{UNet}_{F}$. For the ODOC and Polyp tasks, as expected, the performance of ProCNS increases as the proportion of full annotations increases. The segmentation performance respectively reaches its peak DSC of 89.74% and 86.64% at the 100% full annotation. For the FAZ task, it is notable that at approximately a 30% full annotation, ProCNS reaches a peak DSC of 95.72%. Further increasing the proportion leads to slight declines in model performance. This suggests that a higher proportion of full annotations does not necessarily lead to better performance. Such phenomenon may be attributed to the high sensitivity of sample-wise prototypes within ProCNS and the inherent subjectivity (or noise) presented in full annotations, especially for regions of interest with ambiguous boundaries. Therefore, as the proportion increases, some prototypes might be affected by noise and deviate from accuracy, leading to a decline in model performance. Nonetheless, these results still indicate that providing more annotations to ProCNS can lead to additional performance improvements.
Figure 6: Quantitative analysis on noise suppression concerning the DSC between denoised labels and the corresponding full ground truth. The baseline denotes the DSC between their initial pseudo-labels and their corresponding full ground truth.
@@ 4.7 Effectiveness Analysis on Noise Suppression
The ablation studies in 4.3.1 show that ANPM plays a crucial role in achieving outstanding performance. This is credited to ANPM’s adeptness in iteratively identifying and masking noisy regions within the pseudo-labels. In this subsection, we perform the effectiveness analysis on the noise suppression efficacy of ANPM. Concretely, we first evaluate the DSC between denoised labels and their corresponding full ground truth at various training epochs. As illustrated in Fig. 6, with an increase in epochs, there are consistent increases in DSC across all three tasks. This affirms ANPM’s effectiveness in accurately identifying and suppressing noisy regions. To further validate the role of ANPM, we visualize the masked noisy regions alongside their respective soft pseudo-labels for the three tasks.
As shown in Fig. 7, for the FAZ and polyp tasks, noisy regions predominantly reside at the boundaries of the segmentation targets. As for the ODOC task, a substantial amount of noise is presented at the OD and OC intersection boundary. The soft pseudo-labels generated for these regions through the reassignment strategy are of high quality (as depicted in the red boxes of Fig. 7). This reaffirms the significant role of ANPM. It also elucidates the motivation behind utilizing these unique pseudo-labels for additional supervision over noisy regions.
@@ 5.2 Prototype Calibration
Prototype learning has been widely investigated in segmentation tasks. Conventional prototype-based segmentation algorithms [45, 46, 47] predominantly utilize a MAP operation to generate support prototypes for a specific class, which are subsequently employed to predict the segmentation masks. Recently, within the domain of few-shot segmentation, several studies have innovatively concentrated on prototype calibration and refinement, significantly enhancing model performance. [48] design a Region-enhanced Prototypical Transformer (RPT) to generate multiple prototypes for various regions of a target, integrating them to obtain the ideal prototype; [45] generate extra boundary prototypes for foreground and background by applying automated processes such as erosion and dilation to the ground truth. Nevertheless, the aforementioned approaches come with their drawbacks: the former inevitably increases the computational burden; the latter incorporates extra hyperparameters via the automated algorithms, which make the model more time-consuming to optimize and tune. In contrast, our prototype calibration strategy leverages the model’s potential to progressively detect noisy regions within the mask (initial pseudo-labels) and then employs MAP to secure accurate prototypes, without notably increasing the computational burden or the model’s complexity. It has been demonstrated to be effective in Fig. 6 and Fig. 7.
@@
We expand the exploration into prototype granularity and scale in Table 3. Notably, to our knowledge, this is the first attempt within the realm of WSS to utilize progressive prototype calibration and noise suppression to tackle the issue of the deficiency of prototype semantic representativeness and richness.
5.3 Impacts, Future Works and Limitations
@@
This work holds the promise of facilitating many relevant studies aimed at addressing label-efficient through pseudo-proposal and prototype calibration, encompassing but not limited to weakly supervised learning, semi-supervised learning, noise learning and few-shot learning. Notably, as depicted in Table 5, we have validated that ProCNS can serve as a seamless integration plugin to help other WSS methods break through performance bottlenecks. Furthermore, as shown in Table 6, the reported results do not yield the expected intuitive conclusion that a higher proportion of full annotations results in superior performance. We think that this may be related to the quantity of training data and the inevitable subjective noise in ambiguous boundaries of manual full annotations. Nonetheless, the result that ProCNS outperforms the model employing solely the CE loss when trained under 100% full annotation demonstrates that ProCNS is capable of mitigating the aforementioned issues to a certain extent. However, in the domain of WSS, how to effectively and efficiently locate the most hard-yet-informative regions with inexact annotation data remains an open problem in our research.
Concerning our method’s limitations, ProCNS discerns hard-yet-informative regions based on the degree of prediction volatility for ambiguous regions by the model itself. This means it heavily relies on the model’s potential to identify such regions. If the model lacks sufficient discriminative strength for a particular task, the identified informative regions may be meaningless and even exacerbate the model’s bias. In this case, enhancing the model’s fundamental discriminability might be necessary, for instance, by selecting network architectures or loss functions suited to the task.
@@ 6 Conclusion
This paper proposes a novel WSS framework for medical image segmentation, named ProCNS, which can effectively alleviate model degradation caused by representation bias and noise accumulation. Extensive experiments are conducted on FAZ, ODOC and polyp segmentation tasks, with the superiority of our proposed framework being successfully established.

### v3

@@ 4.7 Analysis on the Impact of Annotation Sparsity
[…]
@@
To explore the impact of the annotation sparsity level on ProCNS and inspired by Zhang et al. [7], we report the results of ProCNS obtained from training samples with five different sparse-versus-full annotated proportions. As shown in Table 10, ProCNS demonstrates superior performance across all five proportions, all of which surpass the performance of $\text{UNet}_{F}$. For the ODOC and Polyp tasks, as expected, the performance of ProCNS increases as the proportion of full annotations increases. The segmentation performance respectively reaches its peak DSC of 89.74% and 86.64% at the 100% full annotation. For the FAZ task, it is notable that at approximately a 30% full annotation, ProCNS reaches a peak DSC of 95.72%. Further increasing the proportion leads to slight declines in model performance. This suggests that higher proportions of full annotations do not necessarily lead to better performance. Such phenomenon may be attributed to the high sensitivity of sample-wise prototypes within ProCNS and the inherent subjectivity (or noise) presented in full annotations, especially for regions of interest with ambiguous boundaries. Therefore, as the proportion increases, some prototypes might be affected by noise and deviate from accuracy, leading to a decline in model performance. Nonetheless, these results still indicate that providing more annotations to ProCNS can lead to additional performance improvements.
Figure 8: Quantitative analysis on noise suppression concerning the DSC between the denoised labels and corresponding full ground truth. The baseline denotes the DSC between their initial pseudo-labels and corresponding full ground truth.
@@ 4.8 Effectiveness Analysis on Noise Suppression
The ablation studies in 4.3.1 show that ANPM plays a crucial role in achieving outstanding performance. This is credited to ANPM’s adeptness in iteratively identifying and masking noisy regions within the pseudo-labels. In this subsection, we perform the effectiveness analysis on the noise suppression efficacy of ANPM. Concretely, we first evaluate the DSC between denoised labels and their corresponding full ground truth at various training epochs. As illustrated in Fig. 8, with an increase in epochs, there are consistent increases in DSC across all three tasks. This affirms ANPM’s effectiveness in accurately identifying and suppressing noisy regions. To further validate the role of ANPM, we visualize the masked noisy regions alongside their respective soft pseudo-labels for the three tasks. As shown in Fig. 9, for the FAZ and polyp tasks, noisy regions predominantly reside at the boundaries of the segmentation targets. As for the ODOC task, a substantial amount of noise is presented at the OD and OC intersection boundary. The soft pseudo-labels generated for these regions through the reassignment strategy are of high quality (as depicted in the red boxes of Fig. 9). This reaffirms the significant role of ANPM. It also elucidates the motivation behind utilizing these unique pseudo-labels for additional supervision over noisy regions.
@@ 5.2 Prototype Calibration
Prototype learning has been widely investigated in segmentation tasks. Conventional prototype-based segmentation algorithms [66, 67, 68] primarily utilize a MAP operation to generate support prototypes for a specific class of interest, which are subsequently employed to predict the segmentation masks. Recently, within the domain of few-shot segmentation, several studies have innovatively explored prototype calibration and refinement, significantly enhancing model performance. Zhu et al. [69] design a Region-enhanced Prototypical Transformer (RPT) to generate multiple prototypes for various regions within the target, integrating them to obtain an ideal prototype; Cheng et al. [66] generate extra boundary prototypes for foreground and background by applying automated processes, such as erosion and dilation, on the ground truth. Nevertheless, the aforementioned approaches come with their drawbacks: the former inevitably increases the computational burden; the latter introduces extra hyperparameters via the automated algorithms, which makes it more time-consuming to optimize and tune the model. In contrast, our prototype calibration strategy leverages the model’s potential to progressively detect noisy regions within the mask (the initial pseudo-labels) and then employs MAP to secure accurate prototypes, without notably increasing the computational burden or the model’s complexity. The effectiveness has been established in Fig. 8 and Fig. 9.
@@
We expand our exploration of prototype granularity and scale in Table 4. To the best of our knowledge, this is the first attempt within the realm of WSS to utilize progressive prototype calibration and noise suppression to tackle the deficiency of prototype semantic representativeness and richness.
5.3 Impacts and Future Works
@@
This work holds promise of facilitating many relevant studies aiming at label-efficient through pseudo-proposal and prototype calibration, encompassing but not limited to weakly supervised learning, semi-supervised learning, noise learning and few-shot learning. Notably, as depicted in Table 8, ProCNS can serve as a seamless integration plugin to help other WSS methods break through performance bottlenecks. Furthermore, as shown in Table 10, a higher proportion of full annotations does not necessarily guarantee better performance, which seems counter-intuitive. We conjecture this might be related to the quantity of the training data and the inevitable subjective noise in the ambiguous boundaries of the manual full annotations. Nonetheless, the result that ProCNS outperforms the model solely employing the CE loss when trained under 100% full annotation demonstrates that ProCNS is capable of mitigating the aforementioned issue to a certain degree. However, effectively and efficiently classifying the most hard-yet-informative regions with inexact annotations remains an open question in our study. Foundational segmentation models pretrained on a large-scale dataset, e.g., SAM [61] and SAM-Med2D [60], demonstrate impressive zero-shot segmentation capabilities and superior performance without requiring additional training, which holds promise in addressing the aforementioned issues (as delineated in Table 9). Looking forward, reasonably leveraging predictions from foundation models to select high-informative regions for fine-grained supervision represents a promising research direction.
Figure 10: Visualization of several failure cases. The overlaid orange denotes masked noisy regions. The obvious prediction errors are highlighted with red boxes.
5.4 Limitations
Concerning our method’s limitation, ProCNS discerns hard-yet-informative regions based on the degree of prediction volatility for ambiguous regions by the model itself. This means it heavily relies on the model’s potential to identify such regions. If the model lacks sufficient discriminative strength for a particular task, the identified informative regions may be meaningless and may even exacerbate the model’s bias. To better illustrate the aforementioned limitation, we train ProCNS employing only the pCE loss at the Initialization stage to diminish the discriminative strength of the preliminary model and then visualize some failure cases. As shown in Fig. 10, ProCNS, like other WSS methods, makes false positive errors for some difficult-to-segment images. The first case demonstrates that when the preliminary model overfits certain non-target regions, the ANPM module identifies a reduced number of noisy regions. The denoised pseudo-labels contain over-confident errors, leading to cumulative model errors. The second case indicates that when the preliminary model exhibits a high degree of prediction volatility, the ANPM module identifies noisy regions much larger than the reliable regions. Consequently, the denoised pseudo-labels are severely eroded, significantly diminishing the semantic richness and accuracy of the calibrated prototypes. This leads to performance degradation. In these two cases, enhancing model’s fundamental discriminability might be necessary, for instance, by employing more advanced network architectures or more appropriate loss functions. The aforementioned conjecture is further evidenced by the fact, as shown in Table 3, that all results without the initial PRSA loss exhibit significant performance gaps compared to those with it, particularly in the relatively challenging polyp lesion segmentation task.
Concerning our experiments’ limitation, given the constraints on the computational resources and the inclusion of multiple tasks and scenarios, all ProCNS-related experiments employ the vanilla UNet rather than more advanced networks. This decision is supported by previous studies [5, 70] indicating that in the field of medical image segmentation, the vanilla UNet does not exhibit significant performance gaps compared to other advanced networks.
@@ 6 Conclusion
This paper proposes a novel WSS framework for medical image segmentation, named ProCNS, which can effectively alleviate model degradation caused by representation bias and noise accumulation. Extensive experiments are conducted on FAZ, ODOC, polyp, nuclei, cardiac multi-structures and whole brain tumor segmentation tasks, with the superiority of our proposed framework being successfully established.